OpenStudy (anonymous):

Solve the system -3y=-2x-7 -4x=-6y-10

OpenStudy (jim766):

first I would get these both in standard form -3y = -2x -7 -4x = -6y -10 +2x +2x +6y +6y 2x -3y = -7 -4x + 6y = -10 now stack em up 2x - 3y = -7 you with me? -4x + 6y = -10

OpenStudy (jim766):

we can eliminate the x's by changing the coefficient from 2 to 4 by multiplying the equation times 2 2(2x - 3y = -7 ) = 4x - 6y = -14 now add the straight down 4x - 6y = -14 -4x + 6y = -10 -------------- 0 + 0 = -28 0 = -28 if both variables are eliminated and the statement is true then the answer is : infinite solutions, the lines are collinear false then the answer is: now solution, the lines are parallel in this case 0 does not equal -28, so there are no solutions any questions?